Minutes of the Tracker Meeting

October 23, 2002


Plyform completed only 2 sandwich panels.  Ricardo will install closeouts in them tomorrow.  Next week they will change to making light trays and will make 3 trays.  Plyform received drawings and materials for the top tray.  Sandro will ask them to do a set of closeouts in aluminum first, as there is no extra carbon-carbon material.  In Italy there are 4 honeycombs to make top/bottom trays.  Sandro suggested that they could make a top tray with the aluminum closeouts for use on the bottom of the live tower. 


The Kapton assembly tool still being measured at Pisa.  The tungsten tooling still is being revised to fix two reference holes.


G&A plans to start next week building new ladders.  Some will be ready for test at the end of next week.


G&A received the CV2500 glue.  The only problem with it is its flamability.  Otherwise it appears not to be hard to use.  They are waiting for the CV15-2500 (which is not flammable).  It was shipped last Friday from SLAC.


Nothing is presently going on at Mipot.  They are waiting for the right set of glues before making more prototype ladders.


G&A promised that next week they will do the first wirebonding tests on the mockup MCMs.


Teledyne is doing the final process evaluation for the MCMs.  There is some concern that the glop-top epoxy has a 150C cure, which can cause warping of boards.  The die attach glue cures at 80C.  They will try the same temperature for the glop top.  They will also look at CV15-2500.  Robert asked if this would give enough protection.  He recalled that the BTEM MCMs were encapsulated with a silicone material, and some wire bonds in the region of the cables were eventually damaged by handling.  Teledyne will bring two MCMs to SLAC next week.  There is still some concern about air bubbles under the pitch adapter that may cause wire bonding problems at G&A.


Data Circuit Systems is rerunning the MCMs.  Only 6 delivered out of the last batch because of a processing problem.


Robert said that the problem with the oscillation of some of the G chips has been more-or-less understood, and a fix was found and successfully tested by FIB.


The EM GTRC chips are diced and in hand.  The tested and diced GTFE chips will come back on Tuesday fully sorted.  They can then be shipped directly to Teledyne.


The mini-MCM was received in Pisa, but the test system has not yet arrived.


Luca is arriving next Monday at 6:15 at SFO.  Tom can pick him up.  He should call Tom's cell phone upon arrival.  Luca said that he can do some work with Mutsumi following his trip to Hawaii, or at SLAC next week.  He recommends to send software for the test system to the student in Pisa asap.


The contract for bottom-tray work finally went to Hytec.  The titanium gusset design and test plan are in place.


Ossie reported that a physicist is arriving at SLAC to make measurements on rectangular invar frames assembled by Ossie.  Ossie made a plywood substrate to support them.  Results can be had by the end of the week, hopefully. 

Ossie reported that heat treating the frames caused the material to deform.  The deformation was enough to break welds in some cases.  It is a very springy material.  Also, it was rusted when it came back from the heat-treat vendor.  Plating will probably therefore be necessary.  Erik will talk to the invar specialty shop about these issues.


Sandro said that customs is holding up shipment of ladders.  Paperwork was not renewed after one year for the temporary importation of the SSDs.


Sandro needs to know if we can we use old face-sheet material for flight production?  How can they be requalified?  What parameters need to be checked?  Tom said that he sent the questions to Erik.  They are stored in a freezer at Plyform.  The expiration date was 6 months after purchase.  Tom will call Fred Gross at GSFC about this.


The top tray assembly tool is at SLAC.  There is a conflict in the CMM area; there is no operator.  Ossie is busy with invar testing.  The insert tool should arrive today.


Erik said that Franz noticed an error in a calculation and now says that the tower-to-tower deflection requirements requires a 190 Hz lateral resonance frequency for the tower.  This emphasises a need to review the random vibe spec.  We need to know what spectrum will be used in LAT testing and whether notching will be done.  Erik will try putting corner gussets and angle braces into the model.  We discussed using the thicker 6-ply face sheets and the heavy core in the bottom tray.